# AirSLAM: An Efficient and Illumination-Robust Point-Line Visual SLAM System

### v2

@@ Abstract
In this paper, we present an efficient visual SLAM system designed to tackle both short-term and long-term illumination challenges. Our system adopts a hybrid approach that combines deep learning techniques for feature detection and matching with traditional backend optimization methods. Specifically, we propose a unified convolutional neural network (CNN) that simultaneously extracts keypoints and structural lines. These features are then associated, matched, triangulated, and optimized in a coupled manner. Additionally, we introduce a lightweight relocalization pipeline that reuses the built map, where keypoints, lines, and a structure graph are used to match the query frame with the map. To enhance the applicability of the proposed system to real-world robots, we deploy and accelerate the feature detection and matching networks using C++ and NVIDIA TensorRT. Extensive experiments conducted on various datasets demonstrate that our system outperforms other state-of-the-art visual SLAM systems in illumination-challenging environments. Efficiency evaluations show that our system can run at a rate of $73\mathrm{Hz}$ on a PC and $40\mathrm{Hz}$ on an embedded platform.
@@ I Introduction
Visual simultaneous localization and mapping (vSLAM) is essential for robot navigation due to its favorable balance between cost and accuracy [1]. Compared to LiDAR SLAM, vSLAM utilizes more cost-effective and compact sensors to achieve accurate localization, thus broadening its range of potential applications. Moreover, cameras can capture richer and more detailed information, which enhances their potential for providing robust localization.
@@
  • We have developed a unified model for both keypoint and line detection, which we call PLNet. To our knowledge, PLNet is the first model capable of simultaneously detecting both point and line features. Furthermore, we associate these two types of features and jointly utilize them for tracking, mapping, and relocalization tasks.
@@
In addition, our engineering contributions include deploying and accelerating feature detection and matching networks using C++ and NVIDIA TensorRT, facilitating their deployment on real robots. We release all the C++ source code at https://github.com/sair-lab/AirSLAM to benefit the community.
@@
  • We extend our system to support both stereo data and stereo-inertial data, increasing its reliability when an inertial measurement unit (IMU) is available.
@@
The remainder of this article is organized as follows. In Section II, we discuss the relevant literature. In Section III, we give an overview of the complete system pipeline. The proposed PLNet is presented in Section IV. In Section V, we introduce the visual-inertial odometry based on PLNet. In Section VI, we present how to optimize the map offline and reuse it online. The detailed experimental results are presented in Section VII to verify the efficiency, accuracy, and robustness of AirSLAM. This article is concluded with limitations in Section VIII.
@@ IV-A Motivation
However, achieving a unified model for keypoint and line detection is challenging, as these tasks typically require different real-image datasets and training procedures. Keypoint detection models are generally trained on large datasets comprising diverse images and depend on either a boosting step or the correspondences of image pairs for training [23, 39, 38]. For line detection, we find wireframe parsing methods [66, 50] can provide stronger geometric cues than the self-supervised models [48, 67] as they are able to detect longer and more complete lines, however, these methods are trained on the Wireframe dataset [68], which is limited in size with only 5,462 discontinuous images. In the following sections, we will address this challenge and demonstrate how to train a unified model capable of performing both tasks. It is important to note that in this paper, the term “line detection” refers specifically to the wireframe parsing task.
@@ IV-B Architecture Design
Line Module: This module takes $\frac{H}{4}\times\frac{W}{4}\times 128$ feature maps as inputs. It consists of a U-Net-like CNN and the line detection header. We modify the U-Net [69] to make it contain fewer convolutional layers and thus be more efficient. The U-Net-like CNN is to increase the receptive field as detecting lines requires a larger receptive field than detecting keypoints. The EPD LOIAlign [50] is used to process the outputs of the line module and finally outputs junctions and lines.
@@ IV-C Network Training
Line Encoding: We adopt the attraction region field [50] to encode line segments. For a line segment $\mathbf{l}=(\mathbf{x_{1}},\mathbf{x_{2}})$, where $\mathbf{x_{1}}$ and $\mathbf{x_{2}}$ are two endpoints of $\mathbf{l}$, and a point $\mathbf{p}$ in the attraction region of $\mathbf{l}$, four parameters and $\mathbf{p}$ are used to encode $\mathbf{l}$:
@@
where $d=\lvert\mathbf{po}\rvert$ and $\mathbf{o}$ is the foot of the perpendicular. $\theta$ is the angle between $\mathbf{l}$ and the $\mathbf{Y}$-axis of the image. $\theta_{1}$ is the angle between between $\mathbf{px_{1}}$ and $\mathbf{po}$. $\theta_{2}$ is the angle between $\mathbf{px_{2}}$ and $\mathbf{po}$. The network can predict these four parameters for point $\mathbf{p}$ and then $\mathbf{l}$ can be decoded through:
@@ V-B Feature Matching
Therefore, to address both the effectiveness problem and efficiency problem, we design a fast and robust line-matching method for illumination-challenging conditions. First, we associate keypoints with line segments through their distances. Assume that $M$ keypoints and $N$ line segments are detected on the image, where each keypoint is denoted as $\mathbf{p}_{i}=(x_{i},y_{i})$ and each line segment is denoted as $\mathbf{l}_{j}=(A_{j},B_{j},C_{j},x_{j,1},y_{j,1},x_{j,2},y_{j,2})$, where $(A_{j},B_{j},C_{j})$ are line parameters of $\mathbf{l}_{j}$ and $(x_{j,1},y_{j,1},x_{j,2},y_{j,2})$ are the endpoints. We first compute the distance between $\mathbf{p}_{i}$ and $\mathbf{l}_{j}$ through:
@@
In this part, we will introduce our 3D feature processing methods, including 3D feature representation, triangulation, i.e., constructing 3D features from 2D features, and re-projection, i.e., projecting 3D features to the image plane. We skip the details of 3D point processing in our system as they are easy to do and similar to other point-based VO and SLAM systems. On the contrary, compared with 3D points, 3D lines have more degrees of freedom, and they are easier to degenerate when being triangulated. Therefore, the 3D line processing will be illustrated in detail.
@@
$\mathbf{L}=\left[\mathbf{n}\mid\mathbf{v}\right]=\underbrace{\left[\begin{array}[]{ccc}\frac{\mathbf{n}}{\lVert\mathbf{n}\lVert}&\frac{\mathbf{v}}{\lVert\mathbf{v}\lVert}&\frac{\mathbf{n}\times\mathbf{v}}{\lVert\mathbf{n}\times\mathbf{v}\lVert}\end{array}\right]}_{\mathbf{U}\in SO(3)}\underbrace{\left[\begin{array}[]{cc}\lVert\mathbf{n}\lVert&\mathbf{0}\\ \mathbf{0}&\lVert\mathbf{v}\lVert\\ \mathbf{0}&\mathbf{0}\end{array}\right]}_{\Sigma_{3\times 2}},$ (7)
@@
$\mathbf{W}=\frac{1}{\sqrt{{\lVert\mathbf{n}\lVert}^{2}+{\lVert\mathbf{v}\lVert}^{2}}}\left[\begin{array}[]{cc}{\lVert\mathbf{n}\lVert}&-{\lVert\mathbf{v}\lVert}\\ {\lVert\mathbf{v}\lVert}&{\lVert\mathbf{n}\lVert}\end{array}\right]\in SO(2).$ (8)
@@ V-C3 Re-projection
where $\mathbf{L}_{c}$ and $\mathbf{L}_{w}$ are Plücker coordinates of 3D line in the camera frame and world frame, respectively. ${\mathbf{R}_{cw}}\in SO(3)$ is the rotation matrix from world frame to camera frame and ${\mathbf{t}_{cw}}\in\mathbb{R}^{3}$ is the translation vector. $\left[\cdot\right]_{\times}$ denotes the skew-symmetric matrix of a vector and ${\mathbf{H}_{cw}}$ is the transformation matrix of 3D lines from world frame to camera frame.
@@ V-F Initial Map
As described in Section III, our map is optimized offline. Therefore, keyframes, map points, and 3D lines will be saved to the disk for subsequent optimization when the visual odometry is finished. For each keyframe, we save its index, pose, keypoints, keypoint descriptors, line features, and junctions. The correspondences between 2D features and 3D features are also recorded. To make the map faster to save, load, and transfer across different devices, the above information is stored in binary form, which also makes the initial map much smaller than the raw data. For example, on the OIVIO dataset [77], our initial map size is only about 2% of the raw data size.
@@ VI-A3 Global Bundle Adjustment
We perform the global bundle adjustment (GBA) after merging duplicated features. The residuals and cost function are similar to Section V-E while the difference is that all keyframes and features will be optimized in this module. In the initial stage of optimization, the re-projection errors of merged features are relatively large due to the VO drift error, so we first iterate 50 times without outlier rejection to optimize the variables to a good rough position, and then iterate another 40 times with outlier rejection.
@@
$\mathbf{E}_{pgo}=\sum{\lVert\text{log}\left(\Delta\tilde{\mathbf{T}}_{ij}^{-1}\mathbf{T}_{i}^{-1}\mathbf{T}_{j}\right)^{\vee}\lVert^{2}_{\Sigma_{ij}}},$ (17)
@@
The systems with online loop detection usually perform the GBA after detecting a new loop, so they undergo multiple repeated GBAs when a scene contains many loops. In contrast, our offline map optimization module only does the GBA after all loop closures are detected, allowing us to reduce the optimization iterations significantly compared with them.
@@ VI-A5 Optimized Map
we save the optimized map for the subsequent map reuse. Compared with the initial map in Section V-F, more information is saved such as the bag of words for each keyframe, the global co-visibility graph, and the scene-dependent junction vocabulary. In the meantime, the number of 3D features has decreased due to the fusion of duplicate map points and 3D lines. Therefore, the optimized map occupies a similar memory to the initial map.
@@ VII Experiments
We use two platforms in the experiments. Most evaluations are conducted on a personal computer with an Intel i9-13900 CPU and a NVIDIA GeForce RTX 4070 GPU. In the efficiency experiment in Section VII-F, we also deploy AirSLAM on an NVIDIA Jetson Orin to prove that our system can achieve good accuracy and efficiency on the embedded platform.
@@ VII-A2 Evaluation Metrics
VII-A3 Results and Analysis
We present the results in Table I. The top-performing results are distinctly highlighted and underlined in order. It can be seen that our PLNet achieves the second-best performance on the Wireframe dataset and the best performance on the YorUrban dataset. On the Wireframe dataset, HAWPv2, the best method, only outperforms our PLNet by 0.5, 0.5, and 0.4 points in $\text{sAP}^{5}$, $\text{sAP}^{10}$, and $\text{sAP}^{15}$, respectively. On the YorUrban dataset, our method surpasses the second-best method by 0.4, 0.8, and 0.9 points on these three metrics, respectively. Overall, we can conclude that our PLNet achieves comparable accuracy with SOTA methods.
Generalizability Analysis
We can also conclude that the generalizability of our PLNet is better than other methods. This conclusion is based on two comparative results between our method and HAWPv2, which is the current best wireframe parsing method. First, on the Wireframe dataset, which also serves as the training dataset, HAWPv2 outperforms our PLNet. However, on the YorUrban dataset, it is surpassed by our method. Second, the previous methods are all evaluated with color inputs in their original paper. Considering that grayscale images are also widely used in vSLAM systems, we train our PLNet with grayscale inputs. We also retrain HAWPv2 and evaluate it using grayscale images for comparison. The result shows that our PLNet significantly outperforms HAWPv2 on both datasets when the inputs are grayscale images. We think the better generalizability comes from our backbone. Other methods are trained on only 5,000 images of the Wireframe dataset, while our backbone is trained on a large diverse dataset, which gives it a stronger feature extraction capability.
Efficiency Analysis
@@ VII-A3 Results and Analysis
  • The FPS of our PLNet is the speed of detecting both keypoints and lines.
@@
Without Loop VINS-Fusion [3] 0.163 0.178 0.316 0.331 0.175 0.102 0.099 0.112 0.110 0.124 0.252 0.178
@@
VII-B Mapping Accuracy
The comparison results are presented in Table II. We evaluate the systems with and without loop detection on 11 sequences. For the comparison without loop detection, our method outperforms other SOTA VIO methods: we achieve the best results on 8 out of 11 sequences. The average translational error of AirVIO is 20% lower than the second-best system, i.e., Kimera-VIO. For the comparison with loop detection, our system achieves comparable performance with ORB-SLAM3 and surpasses other methods. Our AirSLAM achieves the best results on 7 sequences and ORB-SLAM3 achieves the best results on the other 5 sequences, while our average error is a little better than ORB-SLAM3. Another conclusion that can be drawn from Table II is that loop detection significantly improves the accuracy of our system. The average error of our system decreases by 74% after the loop detection.
VII-C Mapping Robustness
@@
VII-C1 Onboard Illumination
VII-C2 Dynamic Illumination
The UMA-VI dataset is a visual-inertial dataset gathered in challenging scenarios with handheld custom sensors. We selected sequences with illumination changes to evaluate our system. As shown in Fig. 8, it contains many sub-sequences where the image suddenly goes dark due to turning off the lights. It is more challenging than the OIVIO dataset for vSLAM systems. As the ground-truth poses are only available at the beginning and the end of each sequence, we disabled the loop closure part from all the evaluated methods.

### v4

@@ Abstract
In this paper, we present an efficient visual SLAM system designed to tackle both short-term and long-term illumination challenges. Our system adopts a hybrid approach that combines deep learning techniques for feature detection and matching with traditional backend optimization methods. Specifically, we propose a unified convolutional neural network (CNN) that simultaneously extracts keypoints and structural lines. These features are then associated, matched, triangulated, and optimized in a coupled manner. Additionally, we introduce a lightweight relocalization pipeline that reuses the built map, where keypoints, lines, and a structure graph are used to match the query frame with the map. To enhance the applicability of the proposed system to real-world robots, we deploy and accelerate the feature detection and matching networks using C++ and NVIDIA TensorRT. Extensive experiments conducted on various datasets demonstrate that our system outperforms other state-of-the-art visual SLAM systems in illumination-challenging environments. Efficiency evaluations show that our system can run at a rate of $73\mathrm{Hz}$ on a PC and $40\mathrm{Hz}$ on an embedded platform. Our implementation is open-sourced: https://github.com/sair-lab/AirSLAM.
@@ I Introduction
Visual simultaneous localization and mapping (vSLAM) is essential for robot navigation due to its favorable balance between cost and accuracy [1]. Compared to LiDAR SLAM, vSLAM utilizes more cost-effective and compact sensors to achieve accurate localization, thus broadening its range of potential applications [2]. Moreover, cameras can capture richer and more detailed information, which enhances their potential for providing robust localization.
@@
  • We develop a unified model for both keypoint and line detection, which we call PLNet. To our knowledge, PLNet is the first model capable of simultaneously detecting both point and line features. Furthermore, we associate these two types of features and jointly utilize them for tracking, mapping, and relocalization tasks.
@@
In addition, our engineering contributions include deploying and accelerating feature detection and matching networks using C++ and NVIDIA TensorRT, facilitating their deployment on real robots. We release all the source code at https://github.com/sair-lab/AirSLAM to benefit the community.
@@
  • We extend our system to support both stereo-only and stereo-inertial data, increasing its reliability when an inertial measurement unit (IMU) is available.
@@
The remainder of this article is organized as follows. In Section II, we discuss the relevant literature. Section III-B presents an overview of the complete system pipeline. The proposed PLNet is presented in Section IV. In Section V, we introduce the visual-inertial odometry based on PLNet. Section VI shows how to optimize the map offline and reuse it online. The detailed experimental results are presented in Section VII to verify the efficiency, accuracy, and robustness of AirSLAM. This article is concluded in Section VIII.
@@ IV-A Motivation
However, achieving a unified model for keypoint and line detection is challenging, as these tasks typically require different real-image datasets and training procedures. Keypoint detection models are generally trained on large datasets comprising diverse images and depend on either a boosting step or the correspondences of image pairs for training [24, 40, 39]. For line detection, we find wireframe parsing methods [67, 51] can provide stronger geometric cues than the self-supervised models [49, 68] as they are able to detect longer and more complete lines. The wireframe includes all prominent straight lines and their junctions within the scene, providing an efficient and accurate representation of large-scale geometry and object shapes [69]. However, these methods are trained on the Wireframe dataset [69], which is limited in size with only 5,462 discontinuous images. In the following sections, we will address this challenge and demonstrate how to train a unified model capable of performing both tasks. It is important to note that in this paper, the term “line detection” refers specifically to the wireframe parsing task.
Figure 4: We use four parameters, $d,\theta,\theta_{1},\theta_{2}$, to encode a line into a point $\mathbf{p}$ within the attraction region field.
@@ IV-B Architecture Design
Line Module: This module takes feature maps from the backbone as inputs. It consists of a U-Net-like CNN and the line detection header. We modify the U-Net [70] to make it contain fewer convolutional layers and thus be more efficient. The U-Net-like CNN is to increase the receptive field as detecting lines requires a larger receptive field than detecting keypoints. The EPD LOIAlign [51] is used to process the outputs of the line module and finally outputs junctions and lines.
@@ IV-C Network Training
Line Encoding: We adopt the attraction region field [51] to encode line segments. As shown in Fig. 4, for a line segment $\mathbf{l}=(\mathbf{x_{1}},\mathbf{x_{2}})$, where $\mathbf{x_{1}}$ and $\mathbf{x_{2}}$ are two endpoints of $\mathbf{l}$, and a point $\mathbf{p}$ in the attraction region of $\mathbf{l}$, four parameters and $\mathbf{p}$ are used to encode $\mathbf{l}$:
@@
where $d$ is the distance from $\mathbf{p}$ to $\mathbf{l}$, $\theta$ is the angle between $\mathbf{l}$ and the $\mathbf{v}$-axis of the image, $\theta_{1}$ is the angle between $\mathbf{px_{1}}$ and the perpendicular line from $\mathbf{p}$ to $\mathbf{l}$, and $\theta_{2}$ is the angle between $\mathbf{px_{2}}$ and the perpendicular line. The network can predict these four parameters for point $\mathbf{p}$ and then $\mathbf{l}$ can be decoded through:
@@ V-B Feature Matching
Therefore, to address both the effectiveness problem and efficiency problem, we design a fast and robust line-matching method for illumination-challenging conditions. First, we associate keypoints with line segments through their distances. Assume that $M$ keypoints and $N$ line segments are detected on the image, where the $i$-th keypoint is denoted as $\mathbf{p}_{i}=(x_{i},y_{i})$ and the $j$-th line segment is denoted as $\mathbf{l}_{j}=(A_{j},B_{j},C_{j},x_{j,1},y_{j,1},x_{j,2},y_{j,2})$, where $(A_{j},B_{j},C_{j})$ are line parameters of $\mathbf{l}_{j}$ and $(x_{j,1},y_{j,1},x_{j,2},y_{j,2})$ are the endpoints. We first compute the distance between $\mathbf{p}_{i}$ and $\mathbf{l}_{j}$ through:
@@
In this part, we will introduce our 3D feature processing methods, including 3D feature representation, triangulation, i.e., constructing 3D features from 2D features, and re-projection, i.e., projecting 3D features to the image plane. For 3D point processing, a 3D point is denoted as $\mathbf{X}\in\mathbb{R}^{3}$. LightGlue [27] is utilized to match keypoints between the left and right images. Successfully matched keypoints are triangulated using stereo disparity information, while unmatched keypoints are triangulated leveraging multi-frame observations. The projection of 3D points onto the image plane is modeled using either the pinhole camera model or the fisheye camera model, depending on the specific camera in use. We skip the details of 3D point processing in our system as they are similar to other point-based VO and SLAM systems [74, 4]. On the contrary, compared with 3D points, 3D lines have more degrees of freedom, and they are easier to degenerate when being triangulated. Therefore, the 3D line processing will be illustrated in detail.
@@
$\mathbf{L}=\left[\mathbf{n}\mid\mathbf{v}\right]=\underbrace{\left[\begin{array}[]{ccc}\frac{\mathbf{n}}{\lVert\mathbf{n}\lVert}&\frac{\mathbf{v}}{\lVert\mathbf{v}\lVert}&\frac{\mathbf{n}\times\mathbf{v}}{\lVert\mathbf{n}\times\mathbf{v}\lVert}\end{array}\right]}_{\mathbf{U}\in\text{SO(3)}}\underbrace{\left[\begin{array}[]{cc}\lVert\mathbf{n}\lVert&\mathbf{0}\\ \mathbf{0}&\lVert\mathbf{v}\lVert\\ \mathbf{0}&\mathbf{0}\end{array}\right]}_{\Sigma_{3\times 2}},$ (7)
@@
[…]
@@ V-C3 Re-projection
[…]
@@ V-F Initial Map
As described in Section III-B, our map is optimized offline. Therefore, keyframes, map points, and 3D lines will be saved to the disk for subsequent optimization when the visual odometry is finished. For each keyframe, we save its index, pose, keypoints, keypoint descriptors, line features, and junctions. The correspondences between 2D features and 3D features are also recorded. To make the map faster to save, load, and transfer across different devices, the above information is stored in binary form, which also makes the initial map much smaller than the raw data. For example, on the OIVIO dataset [79], our initial map size is only about 2% of the raw data size.
@@ VI-A3 Global Bundle Adjustment
We perform the global bundle adjustment (GBA) after merging duplicated features. The form of the residuals and loss functions is similar to that in Section V-E; however, unlike Section V-E, all keyframes and features are jointly optimized, and loop closure residuals are also incorporated into the optimization process. In the initial stage of optimization, the re-projection errors of merged features are relatively large due to the VO drift error, so we first iterate 50 times without outlier rejection to optimize the variables to a good rough position, and then iterate another 40 times with outlier rejection.
@@
$\mathbf{E}_{pgo}=\sum{\lVert\text{Log}\left(\Delta\tilde{\mathbf{T}}_{ij}^{-1}\mathbf{T}_{i}^{-1}\mathbf{T}_{j}\right)\lVert^{2}_{\Sigma_{ij}}},$ (17)
@@
The systems with online loop detection usually perform the GBA after detecting a new loop, so they undergo repeated GBAs when a scene contains many loops. In contrast, our offline map optimization module only does the GBA after all loop closures are detected, allowing us to reduce the optimization iterations significantly compared with them.
@@ VI-A5 Optimized Map
we save the optimized map for subsequent map reuse. Compared with the initial map in Section V-F, more information is saved such as the bag of words for each keyframe, the global co-visibility graph, and the scene-dependent junction vocabulary. In the meantime, the number of 3D features has decreased due to the fusion of duplicate map points and 3D lines. Therefore, the optimized map occupies a similar memory to the initial map.
@@ VII Experiments
We use two platforms in the experiments. Most evaluations are conducted on a personal computer with an Intel i9-13900 CPU and a NVIDIA GeForce RTX 4080 GPU. In the efficiency experiment in Section VII-F, we also deploy AirSLAM on an NVIDIA Jetson Orin to prove that our system can achieve good accuracy and efficiency on the embedded platform.
@@ VII-A2 Evaluation Metrics
  • The FPS of our PLNet is the speed of detecting both keypoints and lines for the Python implementation.
VII-A3 Results and Analysis
We present the results in Table I. The top-performing results are distinctly highlighted and underlined in order. It can be seen that our PLNet achieves the second-best performance on the Wireframe dataset and the best performance on the YorkUrban dataset. On the Wireframe dataset, HAWPv2, the best method, only outperforms our PLNet by 0.5, 0.5, and 0.4 points in $\text{sAP}^{5}$, $\text{sAP}^{10}$, and $\text{sAP}^{15}$, respectively. On the YorkUrban dataset, our method surpasses the second-best method by 0.4, 0.8, and 0.9 points on these three metrics, respectively. Overall, we can conclude that our PLNet achieves comparable accuracy with SOTA methods.
@@ VII-A3 Results and Analysis
Without Loop DROID-SLAM [18] 0.163 0.121 0.242 0.399 0.270 0.103 0.165 0.158 0.102 0.115 0.204 0.186
VINS-Fusion [4] 0.163 0.178 0.316 0.331 0.175 0.102 0.099 0.112 0.110 0.124 0.252 0.178
@@
Generalizability Analysis
We can also conclude that the generalizability of our PLNet is better than other methods. This conclusion is based on two comparative results between our method and HAWPv2, which is the current best wireframe parsing method. First, on the Wireframe dataset, which also serves as the training dataset, HAWPv2 outperforms our PLNet. However, on the YorkUrban dataset, it is surpassed by our method. Second, the previous methods are all evaluated with color inputs in their original paper. Considering that grayscale images are also widely used in vSLAM systems, we train our PLNet with grayscale inputs. We also retrain HAWPv2 and evaluate it using grayscale images for comparison. The result shows that our PLNet significantly outperforms HAWPv2 on both datasets when the inputs are grayscale images. We think the better generalizability comes from our backbone. Other methods are trained on only 5,000 images of the Wireframe dataset, while our backbone is trained on a large diverse dataset, which gives it a stronger feature extraction capability.
Efficiency Analysis
@@ VII-B Mapping Accuracy
The comparison results are presented in Table II. We evaluate the systems with and without loop detection on 11 sequences. For the comparison without loop detection, our method outperforms other VIO methods: we achieve the best results on 8 out of 11 sequences. The average translational error of AirVIO is 20% lower than the second-best system, i.e., Kimera-VIO. For the comparison with loop detection, our system achieves comparable performance with ORB-SLAM3 and DROID-SLAM, which are SOTA traditional and learning-based visual SLAM systems, respectively. Another conclusion that can be drawn from Table II is that loop detection significantly improves the accuracy of our system. The average error of our system decreases by 74% after the loop detection.
VII-C Mapping Robustness